# An Efficient Algorithm for Determining the Equivalence of Zero-one Reaction Networks

### v1

@@ 3.1 Sufficient conditions for vacuous ideals
  • (2)
    for any $i\in\{1,\dots,s\}$, if $P_{i}\neq 0$, then we have $col_{i}(\mathcal{Y})=\vec{0}$,
@@
Let $v=(v_{1},\dots,v_{m})^{\top}$ be defined as in (9). By the condition (1), we have
$P^{\top}v=\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v.$ (11)
Note that $P\neq\vec{0}$. We denote by $P_{i_{1}},\dots,P_{i_{n}}$ all the non-zero coordinates in $P$. So, by (11), we have
$\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v=\sum_{j=1}^{n}P_{i_{j}}v_{i_{j}}.$ (12)
By the condition (2), for any $j\in\{1,\dots,n\}$, we have $col_{i_{j}}(\mathcal{Y})=\vec{0}$. Hence, by (9), for any $j\in\{1,\dots,n\}$,
@@
[…]
@@
A row (column) of a matrix is called a non-zero row (column) if there exists a non-zero element in this row (column). If a row (column) has both positive and negative elements, we say the row (column) changes signs.
@@
For any $d$-dimensional zero-one network $G$ (7), let $f$ be the steady-state system defined as in (8). We denote by $c$ the number of zero columns of the reactant matrix $\mathcal{Y}$. Assume that all the rows of the stoichiometric matrix $\mathcal{N}$ are non-zero and change signs. Then, we have the following statements.
  1. 1.
@@
  2. 2.
    If $c=m-1$ and $d\geq 2$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
  3. 3.
    If $c=m-2$ and $d=s\geq 3$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
  4. 4.
@@
    Recall that any stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ are defined as in (7). By the definition of zero-one network, for any $(i,j)\in\{1,\dots,s\}\times\{1,\dots,m\}$, we have $\beta_{ij}\in\{0,1\}$. So, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 00 or $1$. Hence, if all entries of $\mathcal{Y}$ are 00 (i.e., $c=m$), then all entries of $\mathcal{N}$ are 00 or $1$. It is contrary to the assumption that all the rows of $\mathcal{N}$ are non-zero and change signs. Therefore, there is at least one non-zero column in $\mathcal{Y}$ (i.e., $c<m$).
  2. 2.
@@
    Since all the rows of $\mathcal{N}$ are non-zero and change signs, there exists the element $-1$ in every row of $\mathcal{N}$. By the definition of zero-one network, we have $\alpha_{ij}\in\{0,1\}$ and $\beta_{ij}\in\{0,1\}$. Hence, if there exists an $(i,j)$-entry of $\mathcal{N}$ such that the entry is $-1$ (i.e., $\beta_{ij}-\alpha_{ij}=-1$), then $\alpha_{ij}=1$. Hence, there exists the element $1$ in every row of $\mathcal{Y}$. Since the number of zero columns is $m-1$, we can assume that the first $m-1$ columns of $\mathcal{Y}$ are the zero vector. Note that there exists the element $1$ in every row of $\mathcal{Y}$. So, all the coordinates of the the $m$-th column of $\mathcal{Y}$ must be $1$. Then, we have
    $\mathcal{Y}=\begin{pmatrix}0&\cdots&0&1\\ &\ddots&&\vdots\\ 0&\cdots&0&1\end{pmatrix}.$
@@
    By the definition of zero-one network, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 00 or $1$. Hence, every coordinate of the first $m-1$ columns of $\mathcal{N}$ can only be 00 or $1$. Note that there exists $-1$ in every row vector of $\mathcal{N}$. So, all the coordinates of the the $m$-th column of $\mathcal{N}$ are $-1$. Thus, $\mathcal{N}$ has the following form
    $\mathcal{N}=\begin{pmatrix}&&&-1\\ &\mathcal{A}_{1}&&\vdots\\ &&&-1\end{pmatrix},$ (14)
@@
    where all the entries of the $s\times(m-1)$ submatrix $\mathcal{A}_{1}$ are 00 or $1$. Recall that for any $i\in\{1,\dots,s\}$, we denote by $row_{i}(\mathcal{N})$ the $i$-th row of $\mathcal{N}$. Since $d\geq 2$ (note here, ${\tt rank}({\mathcal{N}})=d$), there exist $i,j\in\{1,\dots,s\}$ such that $P^{\top}:=row_{i}(\mathcal{N})-row_{j}(\mathcal{N})\neq\vec{0}^{\top}$. So, if $P_{k}\neq 0$, then by (14), we have $k\in\{1,\dots,m-1\}$. Note that for any $k\in\{1,\dots,m-1\}$, the $k$-th column vector of $\mathcal{Y}$ is a zero vector. Therefore, by Theorem 1, we have $\langle f\rangle=\langle 1\rangle$.
  3. 3.
@@
    Assume that the first $m-2$ columns of $\mathcal{Y}$ are the zero vector. Then, by the definition of zero-one network, we have
@@
    where the entries of the $s\times 2$ submatrix $\mathcal{A}_{2}$ are 00 or $1$. Then, similar to the proof of (2), by the definition of zero-one network, we have
@@
    where the entries of the $s\times(m-2)$ submatrix $\mathcal{A}_{3}$ are 00 or $1$. Since all rows of $\mathcal{N}$ are non-zero and change signs, there exists the element $-1$ in every row of $\mathcal{N}$. Hence, the rows of the $s\times 2$ submatrix $\mathcal{A}_{4}$ must be in the set
@@
      Assume that there exist $k<s$ different row vectors in (16) appearing in $\mathcal{A}_{4}$. Since $k<s$, at least two rows in $\mathcal{A}_{4}$ are the same. Assume that the $i$-th and the $j$-th ($i,j\in\{1,\dots,s\}$) rows in $\mathcal{A}_{4}$ are the same. Define $P^{\top}:=row_{i}(\mathcal{N})-row_{j}(\mathcal{N})$. Then, we have $P_{m-1}=P_{m}=0$. Since ${\tt rank}(\mathcal{N})=d=s$, we have $P\neq\vec{0}$. Therefore, for any $k\in\{1,\dots,m\}$, if $P_{k}\neq 0$, we have $k\in\{1,\dots,m-2\}$. Note that for any $k\in\{1,\dots,m-2\}$, the $k$-th column of $\mathcal{Y}$ is the zero vector. Hence, by Theorem 1, we have $\langle f\rangle=\langle 1\rangle$.
@@
  4. 4.
    Since there is no zero columns in $\mathcal{Y}$, by (9), for any $i\in\{1,\dots,m\}$, $v_{i}=\kappa_{i}\prod_{j=1}^{s}x^{\alpha_{ji}}$, where $\sum\limits_{j=1}^{s}\alpha_{ji}\geq 0$ since $\alpha_{ji}\in\{0,1\}$. Hence, by (8), all terms in $f$ contain the variables $x_{i}$’s. So, in ${\mathbb{Q}}(\kappa)[x]$, $1\notin\langle f\rangle$. Hence, we have $\langle f\rangle\neq\langle 1\rangle$.
@@
(1) Consider the following zero-one network:
[…]
$0\displaystyle 0\xrightarrow{\kappa_{4}}X_{1},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{5}}$ 0 .
@@
Note that there are $4$ zero columns in $\mathcal{Y}$ (i.e., $c=4=m-1$). Hence, by Theorem 2 (2), we have $\langle f\rangle=\langle 1\rangle$. Note that
@@
(2) Consider the following zero-one network:
[…]
@@
Note that there are $3$ zero columns in $\mathcal{Y}$ (i.e., $c=3=m-2$). Hence, by Theorem 2 (3), we have $\langle f\rangle=\langle 1\rangle$. Note that
@@
(3) Consider the following zero-one network:
@@
Note that there is no zero columns in $\mathcal{Y}$ (i.e., $c=0$). Hence, by Theorem 2 (4), we have $\langle f\rangle\neq\langle 1\rangle$. Note that
@@ 3.2 Sufficient conditions for equivalence
Theorem 3.
For any two networks $G$ and $G^{^{\prime}}$, let $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ be the corresponding stoichiometric matrices, and let $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ be the reactant matrices. Let $f$ and $f^{^{\prime}}$ be the steady-state systems defined as in (8). If
  • (1)
    $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, and
  • (2)
    for any $i\in\{1,\dots,s\}$, we have
    $row_{i}(\mathcal{N})\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}^{^{\prime}}),\dots,row_{s}(\mathcal{N}^{^{\prime}})\},$
    and
    $row_{i}(\mathcal{N}^{^{\prime}})\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}),\dots,row_{s}(\mathcal{N})\},$
where $\operatorname{span}_{\mathbb{Q}}\{\cdot\}$ means the rational vector space spanned by a set of rational vectors, then $\langle f\rangle=\langle f^{^{\prime}}\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
@@
Let $v=(v_{1},\dots,v_{m})^{\top}$ and $v^{^{\prime}}=(v_{1}^{^{\prime}},\dots,v_{m}^{^{\prime}})^{\top}$ be defined as in (9) corresponding to $G$ and $G^{\prime}$ respectively, where
$v_{i}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\alpha_{ji}},\;\;v_{i}^{^{\prime}}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\alpha_{ji}^{^{\prime}}}.$ (17)
By $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, for any $i\in\{1,\dots,s\}$ and for any $j\in\{1,\dots,s\}$, $\alpha_{ij}=\alpha_{ij}^{^{\prime}}$. Hence, by (17), we have $v=v^{^{\prime}}$. So, the conclusion follows from (8) and the condition (2). ∎
@@
Consider the following two networks $G$ and $G^{^{\prime}}$.
@@
$G^{^{\prime}}:$ $\quad X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{2}.$
Note that the two networks $G$ and $G^{^{\prime}}$ have the same set of reactants. Hence, the reactant matrices $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ are equal. The stoichiometric matrices $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ are as follows.
$\mathcal{N}=\begin{pmatrix}-1&0&1\\ 0&-1&1\end{pmatrix},\quad\mathcal{N}^{^{\prime}}=\begin{pmatrix}-1&1&0\\ 0&-1&1\end{pmatrix}.$
Note that $row_{1}(\mathcal{N})=row_{1}(\mathcal{N}^{^{\prime}})+row_{2}(\mathcal{N}^{^{\prime}})$, $row_{2}(\mathcal{N})=row_{2}(\mathcal{N}^{^{\prime}})$ and $row_{1}(\mathcal{N}^{^{\prime}})=row_{1}(\mathcal{N})-row_{2}(\mathcal{N})$. So, by Theorem 3, the ideals generated by $f$ and $f^{^{\prime}}$ are equal.
@@
We remark that although Theorem 1 and Theorem 3 are straightforward to prove, the hypotheses in these theorems can not be directly checked by a computer, while Theorem 2 is explicit enough. In the next section, we will solve this issue by Algorithms 2 and 3.

### v2

@@ 3.2 Sufficient conditions for vacuous ideals
  2. (II)
    for any $i\in\{1,\dots,s\}$, $P_{i}\neq 0$ implies $col_{i}(\mathcal{Y})=\vec{0}$,
@@
Let $v=(v_{1},\dots,v_{m})^{\top}$ be defined as in (9). By condition (I), we have $P\neq\vec{0}$. And, we have
$\sum_{j=1}^{n}P_{i_{j}}v_{i_{j}}=P^{\top}v=\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v,$ (12)
where $P_{i_{1}},\dots,P_{i_{n}}$ denotes all the non-zero coordinates in $P$. By condition (II) , for any $j\in\{1,\dots,n\}$, we have $col_{i_{j}}(\mathcal{Y})=\vec{0}$. Hence, by (9), for any $j\in\{1,\dots,n\}$,
@@
$X_{1}\xrightarrow{\kappa_{1}}X_{2},\;\;X_{2}\xrightarrow{\kappa_{2}}X_{1},\;\;0\xrightarrow{\kappa_{3}}X_{2}.$
@@
A row (column) of a matrix is called a non-zero row (column) if there exists a non-zero element in this row (column). If a row (column) has both positive and negative elements, we say the row (column) changes signs. Notice that if there exists one non-zero row of a stoichiometric matrix $\mathcal{N}$ that does not change sign, then the corresponding network admits no positive steady states.
@@
For any $d$-dimensional zero-one network $G$ (7), let $f$ be the steady-state system defined as in (8). We denote by $c$ the number of zero columns of the reactant matrix $\mathcal{Y}$. Assume that all the non-zero rows of the stoichiometric matrix $\mathcal{N}$ change signs. Then, we have the following statements.
  1. (I)
@@
  2. (II)
    If c=m1c=m-1 and $d\geq 2$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
  3. (III)
    If c=m2c=m-2 and $d=s\geq 3$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
  4. (IV)
@@
    Recall that any stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ are defined as in (7). By the definition of zero-one network, for any $(i,j)\in\{1,\dots,s\}\times\{1,\dots,m\}$, we have $\beta_{ij}\in\{0,1\}$. So, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 0 or $1$. Hence, if all entries of $\mathcal{Y}$ are 0 (i.e., $c=m$), then every entry of $\mathcal{N}$ is 0 or $1$. By the definition of zero-one network, $\mathcal{N}$ is not a zero matrix. So, there exists a non-zero row of $\mathcal{N}$ whose entries are 0 or $1$, which is contrary to the assumption that all the non-zero rows of $\mathcal{N}$ change signs. Therefore, there is at least one non-zero column in $\mathcal{Y}$ (i.e., $c<m$).
  2. (II)
@@
    Since all the non-zero rows of $\mathcal{N}$ change signs, there exists the element 1-1 in every non-zero row of $\mathcal{N}$. We define a new matrix $\mathcal{N}^{*}$ by removing the zero rows from $\mathcal{N}$. Also, we define $\mathcal{Y}^{*}$ by removing the same rows from $\mathcal{Y}$. By the definition of zero-one network, we have $\alpha_{ij}\in\{0,1\}$ and $\beta_{ij}\in\{0,1\}$. Hence, if there exists an $(i,j)$-entry of $\mathcal{N}^{*}$ such that the entry is 1-1 (i.e., $\beta_{ij}-\alpha_{ij}=-1$), then $\alpha_{ij}=1$. Hence, there exists the element $1$ in every row of $\mathcal{Y}^{*}$. Since the number of zero columns of $\mathcal{Y}$ is m1m-1, we can assume that the first m1m-1 columns of $\mathcal{Y}$ are the zero vectors. Then, by the definition of $\mathcal{Y}^{*}$, the first m1m-1 columns of $\mathcal{Y}^{*}$ are the zero vectors. Note that there exists the element $1$ in every row of $\mathcal{Y}^{*}$. So, all the coordinates of the the $m$-th column of $\mathcal{Y}^{*}$ must be $1$. Since $d\geq 2$ (note here, ${{\rm rank}}({\mathcal{N}})=d$), the matrix $\mathcal{Y}^{*}$ has more than one row. Then, we have
    ${\mathcal{Y}^{*}}=\begin{pmatrix}0&\cdots&0&1\\ &\ddots&&\vdots\\ 0&\cdots&0&1\end{pmatrix}$.
@@
    By the definition of zero-one network, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}^{*}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 0 or $1$. Hence, every coordinate of the first m1m-1 columns of $\mathcal{N}^{*}$ can only be 0 or $1$. Note that there exists 1-1 in every row vector of $\mathcal{N}^{*}$. So, all the coordinates of the the $m$-th column of $\mathcal{N}^{*}$ are 1-1. Thus, $\mathcal{N}^{*}$ has the following form
    ${\mathcal{N}^{*}}=\begin{pmatrix}&&&-1\\ &\mathcal{A}_{1}&&\vdots\\ &&&-1\end{pmatrix}$, (14)
@@
    where all the entries of the submatrix $\mathcal{A}_{1}$ are 0 or $1$. Recall that for any $i\in\{1,\dots,s\}$, $row_{i}(\mathcal{N})$ denotes the $i$-th row of $\mathcal{N}$. Since $d\geq 2$, there exist two non-zero rows $row_{i}(\mathcal{N})$ and $row_{j}(\mathcal{N})$ such that $P^{\top}:=row_{i}(\mathcal{N})-row_{j}(\mathcal{N})\neq\vec{0}^{\top}$. So, if $P_{k}\neq 0$, then by (14), we have $k\in\{1,\dots,m-1\}$. Note that for any $k\in\{1,\dots,m-1\}$, the $k$-th column vector of $\mathcal{Y}$ is a zero vector. Therefore, by Theorem 2, we have $\langle f\rangle=\langle 1\rangle$.
  3. (III)
@@
    Assume that the first m2m-2 columns of $\mathcal{Y}$ are the zero vector. Then, by the definition of zero-one network, we have
@@
    where the entries of the $s\times 2$ submatrix $\mathcal{A}_{2}$ are 0 or $1$. Since $d=s$, all rows of $\mathcal{N}$ are non-zero. Then, similar to the proof of (2), by the definition of zero-one network, we have
@@
    where the entries of the $s\times(m-2)$ submatrix $\mathcal{A}_{3}$ are 0 or $1$. Since all rows of $\mathcal{N}$ change signs, there exists the element 1-1 in every row of $\mathcal{N}$. Hence, the rows of the $s\times 2$ submatrix $\mathcal{A}_{4}$ must be in the set
@@
      Assume that there exist $k<s$ different row vectors in (16) appearing in $\mathcal{A}_{4}$. Since $k<s$, at least two rows in $\mathcal{A}_{4}$ are the same. Assume that the $i$-th and the $j$-th ($i,j\in\{1,\dots,s\}$) rows in $\mathcal{A}_{4}$ are the same. Define $P^{\top}:=row_{i}(\mathcal{N})-row_{j}(\mathcal{N})$. Then, we have $P_{m-1}=P_{m}=0$. Since ${{\rm rank}}(\mathcal{N})=d=s$, we have $P\neq\vec{0}$. Therefore, for any $k\in\{1,\dots,m\}$, if $P_{k}\neq 0$, we have $k\in\{1,\dots,m-2\}$. Note that for any $k\in\{1,\dots,m-2\}$, the $k$-th column of $\mathcal{Y}$ is the zero vector. Hence, by Theorem 2, we have $\langle f\rangle=\langle 1\rangle$.
@@
  4. (IV)
    Since there is no zero columns in $\mathcal{Y}$, by (9), for any $i\in\{1,\dots,m\}$, $v_{i}=\kappa_{i}\prod_{j=1}^{s}x^{\alpha_{ji}}$, where $\sum\limits_{j=1}^{s}\alpha_{ji}>0$ since $\alpha_{ji}\in\{0,1\}$. Hence, by (8), for any $i\in\{1,\dots,s\}$, if $row_{i}(\mathcal{N})$ is a zero vector, then $f_{i}=0$. By (8), for any $i\in\{1,\dots,s\}$, if $row_{i}(\mathcal{N})$ is non-zero, then all terms in $f_{i}$ contain the variables $x_{j}$’s. So, in ${\mathbb{Q}}(\kappa)[x]$, $1\notin\langle f\rangle$. Hence, we have $\langle f\rangle\neq\langle 1\rangle$.
@@
(I) Consider the following zero-one network:
$0\xrightarrow{\kappa_{1}}X_{3},\;\;0\xrightarrow{\kappa_{2}}X_{2},\;\;0\xrightarrow{\kappa_{3}}X_{2}+X_{3},$
$0\xrightarrow{\kappa_{4}}X_{1},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{5}}0.$
@@
Note that there are $4$ zero columns in $\mathcal{Y}$ (i.e., c=4=m1c=4=m-1). Hence, by Theorem 3 (II), we have $\langle f\rangle=\langle 1\rangle$. Note that
@@
(II) Consider the following zero-one network:
$0\xrightarrow{\kappa_{1}}X_{3},\;\;0\xrightarrow{\kappa_{2}}X_{2},\;\;0\xrightarrow{\kappa_{3}}X_{2}+X_{3},$
@@
Note that there are $3$ zero columns in $\mathcal{Y}$ (i.e., c=3=m2c=3=m-2). Hence, by Theorem 3 (III), we have $\langle f\rangle=\langle 1\rangle$. Note that
@@
(III) Consider the following zero-one network:
@@
Note that there is no zero columns in $\mathcal{Y}$ (i.e., $c=0$). Hence, by Theorem 3 (IV), we have $\langle f\rangle\neq\langle 1\rangle$. Note that
@@
Remark 2.
We remark that although Theorem 2 is straightforward to prove, the hypotheses in the theorems cannot be directly checked by a computer, while Theorem 3 is explicit enough. In Section 4, we solve this issue by Algorithm 3.
@@ 3.3 Sufficient conditions for equivalence
Lemma 1.
For any two networks $G$ and $\hat{G}$ with $s$ species and $m$ reactions, let $\mathcal{N}$ and $\hat{\mathcal{N}}$ be the corresponding stoichiometric matrices, and let $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ be the reactant matrices. Let $f$ and $\hat{f}$ be the steady-state systems defined as in (8). If
  1. (I)
    $\mathcal{Y}=\hat{\mathcal{Y}}$, and
  2. (II)
    for any $i\in\{1,\dots,s\}$, we have row_i(N) ∈span_Q {row_1(^N), …, row_s(^N)}, and row_i(^N) ∈span_Q {row_1(N), …, row_s(N)},
where $\operatorname{span}_{\mathbb{Q}}\{\cdot\}$ means the rational vector space spanned by a set of rational vectors, then $\langle f\rangle=\langle\hat{f}\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
@@
Let $v=(v_{1},\dots,v_{m})^{\top}$ and $\hat{v}=(\hat{v}_{1},\dots,\hat{v}_{m})^{\top}$ be defined as in (9) corresponding to $G$ and $\hat{G}$ respectively, where
$v_{i}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\alpha_{ji}},\;\;\hat{v}_{i}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\hat{\alpha}_{ji}}.$ (17)
By $\mathcal{Y}=\hat{\mathcal{Y}}$, for any $i\in\{1,\dots,s\}$ and for any $j\in\{1,\dots,s\}$, $\alpha_{ij}=\hat{\alpha}_{ij}$. Hence, by (17), we have $v=\hat{v}$. So, the conclusion follows from (8) and condition (II) . ∎
Remark 3.
If $G$ has less reactions than $\hat{G}$ does, then we can check if there exists a subnetwork of $\hat{G}$ such that the subnetwork is equivalent to $G$. Here, by “subnetwork”, we mean the network consists of some reactions of $\hat{G}$ and it has the same stoichiometric subspace as $\hat{G}$, where stoichiometric subspace means the column space of the stoichiometric matrix. By [7, 6], if a subnetwork admits multistability or Hopf bifurcations, then the original network also admits multistability or Hopf bifurcations.
@@
Consider the following two networks $G$ and $\hat{G}$.
@@
$\hat{G}$: $\quad X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{2}.$
Note that the two networks $G$ and $\hat{G}$ have the same set of reactants. Hence, the reactant matrices $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ are equal. The stoichiometric matrices $\mathcal{N}$ and $\hat{\mathcal{N}}$ are as follows.
$\mathcal{N}=\begin{pmatrix}-1&0&1\\ 0&-1&1\end{pmatrix},\quad\hat{\mathcal{N}}=\begin{pmatrix}-1&1&0\\ 0&-1&1\end{pmatrix}$.
Note that
[…]
So, by Lemma 1, the ideals generated by $f$ and $\hat{f}$ are equal.
@@
Notice that each network is uniquely determined by a pair of stoichiometric matrix $\mathcal{N}$ and reactant matrix $\mathcal{Y}$. However, a stoichiometric matrix $\mathcal{N}$ might correspond to many different reactant matrices. For example, if a network has the following stoichiometric matrix
$\mathcal{N}=\begin{pmatrix}1&1&0&-1\\ 1&-1&-1&1\end{pmatrix}$, (18)
then the reactant matrix can be one of the following two matrices
${\mathcal{Y}}_{1}=\begin{pmatrix}0&0&0&1\\ 0&1&1&0\end{pmatrix},\quad{\mathcal{Y}}_{2}=\begin{pmatrix}0&0&1&1\\ 0&1&1&0\end{pmatrix}$.
So, for two networks, if their stoichiometric matrices $\mathcal{N}$ and $\hat{\mathcal{N}}$ have different row spaces, then we do not need to compare the reactant matrices $\mathcal{Y}$ and $\hat{\mathcal{Y}}$. That means in practice, we can first check condition (II) of Lemma 1, and if it holds, we check condition (I). In Section 4, we check if condition (II) holds by checking if the reduced row echelon forms of the matrices $\mathcal{N}$ and $\hat{\mathcal{N}}$ are equal. See Step 4 in Algorithm 1.